# Range Resolution of Near-field MIMO Sensing

### v1

@@ Abstract
The radiative near-field and integration of sensing capabilities are seen as two key components of the next generation of wireless communication systems. In this paper, the sensing performance of a narrowband near-field system is investigated for several practical antenna array geometries and configurations, namely SIMO/MISO and MIMO. In the SIMO/MISO configuration, the antenna aperture is exploited only a single time for either transmit or receive signal processing, while the MIMO configuration exploits both TX and RX processing. Analytical derivations, supported by simulations, show that the MIMO processing improves the maximum near-field range and sensing resolution by approximately a factor of 1.4 as compared to single-aperture systems. The value of the improvement factor is consistent for all considered array geometries. Finally, using a quadratic approximation of the array factor, an analytical improvement factor of $\sqrt{2}$ is derived, clarifying the observed improvements and validating the numerical results.
@@ I-A Problem Statement
The radiative near-field (NF) systems are regarded as a key enabler of the next generation of wireless systems [1]. Their ability to create beamspots in the distance domain enables user multiplexing both in range and angle [2, 3]. This unique property has also attracted the attention of the sensing community as beamfocusing in the range domain can be utilized to localize the targets with limited bandwidth [4, 5]. Despite these advantages, near-field sensing performance is subject to significant shortcomings, such as limited resolution, maximum near-field range and poor peak-to-sidelobe level (PSL) performance. The recent research regarding near-field sensing has largely focused on single-array geometry and single-aperture systems [6, 3, 5, 7, 4]. To improve the near-field sensing, the performance gains from MIMO sensing have to be identified and quantified for different array geometries.
@@ I-B Relevant works
In [6], the near-field beam focusing was derived for a uniform square array, based on electromagnetic principles. Generalizations of the array factor (AF) in the near-field for arbitrary uniform rectangular array (URA) and uniform planar circular array (UPCA) are provided in [3]. This work studies the beamfocusing property from a communication-centric perspective. The near-field communications performance and AF of a uniform circular array (UCA) were discussed in [7].
In [4] the near-field sensing resolution was demonstrated using a testbed equipped with a uniform linear antenna array (ULA). An approximation of the near-field range resolution for the same array architecture and SIMO/MISO case is provided in [5]. The sensing performance of a near-field system for point and extended targets was investigated in [8]. This work discusses the NF performance of the SIMO/MISO and MIMO systems regarding the Cramér–Rao lower bound, but does not directly discuss the gains from MIMO for point targets.
@@ I-C Contributions
In this work, the NF sensing performance is evaluated for several array geometries: ULA, UCA, URA and UPCA. The sensing performance is studied for two cases SIMO/MISO - single aperture and MIMO - two identical exactly collocated apertures. The gains from employing MIMO are assessed in terms of resolution, maximum NF range and PSL level of the ambiguity function. Based on a quadratic approximation, the analytical improvement factor is derived, which matches well with the numerical simulations.
@@ II-A Ambiguity function of near-field sensing system
Consider a collocated narrowband antenna array with $M$ transmitters and $N$ receivers. The location of the of the $m$th transmitter is given by $\mathbf{p}_{m}$ and $n$th receiver by $\mathbf{p}_{n}$. The antenna elements are assumed to be isotropic and mutual coupling is considered negligible. The distance between each antenna element and some point in space is $d_{m}(\mathbf{p})=\left|\left|\mathbf{p}-\mathbf{p}_{m}\right|\right|_{2}$ for transmitters and $d_{n}(\mathbf{p})=\left|\left|\mathbf{p}-\mathbf{p}_{n}\right|\right|_{2}$ for receivers. The bistatic distance is denoted as $d_{m,n}(\mathbf{p})=d_{m}(\mathbf{p})+d_{n}(\mathbf{p})$. The region of interest is located within the radiative near-field of the array and the propagation is modeled as line-of-sight. The channel can be written as
@@
$\Delta d_{m}(\mathbf{p}^{\prime},\mathbf{p})$ $=\Delta d_{m=0}(\mathbf{p}^{\prime},\mathbf{p})+\delta d_{m}(\mathbf{p}^{\prime},\mathbf{p})$ (4)
$\Delta d_{n}(\mathbf{p}^{\prime},\mathbf{p})$ $=\Delta d_{n=0}(\mathbf{p}^{\prime},\mathbf{p})+\delta d_{n}(\mathbf{p}^{\prime},\mathbf{p}).$ (5)
This allows to express the matched filter in terms of array factors as
$\mathcal{A}(\mathbf{p}^{\prime},\mathbf{p})=e^{-j2\pi\frac{c}{f}\Delta d_{m=0}(\mathbf{p}^{\prime},\mathbf{p})}\underbrace{\frac{1}{\sqrt{M}}\sum_{m\in M}e^{-j2\pi\frac{f}{c}\delta d_{m}(\mathbf{p}^{\prime},\mathbf{p})}}_{\mathrm{AF_{TX}}(\mathbf{p}^{\prime},\mathbf{p})}$
×$\qquad\times e^{-j2\pi\frac{c}{f}\Delta d_{n=0}(\mathbf{p}^{\prime},\mathbf{p})}\underbrace{\frac{1}{\sqrt{N}}\sum_{n\in N}e^{-j2\pi\frac{f}{c}\delta d_{n}(\mathbf{p}^{\prime},\mathbf{p})}}_{\mathrm{AF_{RX}}(\mathbf{p}^{\prime},\mathbf{p})}.$ (6)
@@
The MIMO configuration improves the ambiguity function by introducing a multiplication by a second array factor compared to SIMO/MISO case, where only a single array factor defines the ambiguity function. In the near-field region, the AF depends both on angle and distance from the array, unlike in the far-field, where it is only angle dependent. The distance sensitivity results in an array gain that varies with distance and facilitates the beamfocusing. The distance dependence of the AF leverages new applications allowing to discriminate targets and users located at the same angle but different distances without the use of bandwidth, which is not possible in the far-field.
@@ II-B Near-field 3 dB beamdepth
Consider a target and points of interest located along the array normal, this allows to reduce the vectors $\mathbf{p}$, and $\mathbf{p}^{\prime}$ to scalars $d$ and $d^{\prime}$, which represent the distance from the weight center of the array along the normal. The AF in the near-field can be calculated analytically for different array architectures. [6, 3, 7]. Due to the Fresnel approximation of distance in the NF of the array, the argument of the AF shares the same formula for different array architectures
$x(d,d^{\prime})$ $=ad_{\mathrm{FA}}d_{ver}$ (8)
where $a$ is a scalar dependent on the array geometry, $d_{\mathrm{FA}}=2D^{2}/\lambda$ is the Fraunhofer distance of the array, $D$ is the array aperture, $\lambda$ is the wavelength and $d_{ver}=\left|\frac{1}{d}-\frac{1}{d^{\prime}}\right|=\frac{|d^{\prime}-d|}{d^{\prime}d}$ is the absolute value of vergence difference [9]. For brevity in the following, the shorthand $x$ is used to represent $x(d,d^{\prime})$. Consider some near-field ambiguity function given by $|\mathcal{A}(x)|^{2}$. For a SISO/SIMO system, the ambiguity function will be defined by a single array factor of the transmit or receive aperture $|\mathcal{A}(x)|_{\mathrm{SIMO/MISO}}^{2}=|\mathrm{AF}(x)|^{2}$. For the MIMO scheme, the ambiguity function will be a product of two array factors. For simplicity, the transmit and receive arrays are considered to be identical and exactly collocated, resulting in a monostatic system $|\mathcal{A}(x)|_{\mathrm{MIMO}}^{2}=|\mathrm{AF}(x)|^{4}$.
The range resolution is evaluated in terms of the radial half-power beamwidth of the ambiguity function, also referred to as the beamdepth (BD). The 3 dB beamdepth is calculated by computing the argument for which the ambiguity function reaches half of the maximum value. Considering normalized ambiguity function it can be written as $|\mathcal{A}(x_{\mathrm{3dB}})|^{2}=0.5$. The argument value is then used to calculate the distance $d$ at which the $-3$ dB value is reached by solving (8) for $d$ resulting in
@@
Formula (9) gives two distances for which the $-3$ dB value is achieved for target at $d^{\prime}$, where $\alpha=x_{\mathrm{3dB}}/a$. The parameter $a$ depends only on the array geometry, while $x_{\mathrm{3dB}}$ depends on both the array architecture and configuration SIMO/MISO or MIMO. As $d_{\mathrm{3dB}}$ should always be positive, the following constraint arises $d^{\prime}<\frac{d_{\mathrm{FA}}}{\alpha}$. Given the $-3$ dB crossing points, 3 dB beamdepth (BD) is
$BD$ =$=\begin{cases}\frac{2\alpha d_{\mathrm{FA}}d^{\prime 2}}{d_{\mathrm{FA}}^{2}-\alpha^{2}d^{\prime 2}},&d^{\prime}<\frac{d_{\mathrm{FA}}}{\alpha},\\ \infty,&d^{\prime}\geq\frac{d_{\mathrm{FA}}}{\alpha}.\end{cases}$ (10)
For $d$ greater or equal to $\frac{d_{\mathrm{FA}}}{\alpha}$ the beamdepth extends from $d_{\mathrm{FA}}/\alpha$ to infinity. Note that parameter $\alpha$ determines both the 3 dB beamdepth resolution and the maximum near-field distance for which the BD is finite. In the following sections, the $\alpha$ parameter is computed for different array architectures to assess the resolution and maximum near-field sensing range.
@@ III-D2 MIMO
ULA 6.952 4.968 1.399
UCA 5.735 4.151 1.382
URA 9.936 7.072 1.405
UPCA 7.088 5.104 1.389
Table I presents the summarized values of the parameter $\alpha$ for each considered array geometry and processing. By comparing the $\alpha$ values across antenna geometries and processing a few observations can be made. The MIMO processing improves the $\alpha$ parameter approximately by the same factor of $1.4$ regardless of the array geometry. This corresponds to $1.4$-fold improvement in the NF sensing resolution and maximum NF range.
The consistent ratio $\frac{\alpha_{\mathrm{SIMO/MISO}}}{\alpha_{\mathrm{MIMO}}}$ observed across various array geometries can be attributed to quadratic decay of AF within the mainlobe. Based on this, the mainlobe can be accurately approximated as a quadratic function [10]. Since the considered normalized AF is even, with a maximum value of 1, for SIMO/MISO configuration it can be approximated as
$|\mathrm{AF}(x)|^{2}$ $\approx 1-cx^{2},$ (23)
where $c$ is some scalar that controls the decay rate and depends on the array geometry. The $-3$ dB value is obtained for argument $x_{\mathrm{3dB}}^{\mathrm{SIMO/MISO}}=\frac{\sqrt{2}}{2\sqrt{c}}$.
Next, in the MIMO case, the approximation from (23) is squared. To keep the approximation quadratic for both scenarios, higher order terms are neglected, resulting in
$|\mathrm{AF}(x)|^{4}$ $\approx 1-2cx^{2}.$ (24)
The $-3$ dB value is obtained for argument $x_{\mathrm{3dB}}^{\mathrm{MIMO}}=\frac{1}{2\sqrt{c}}$. Based on the quadratic approximation the ratio between the SIMO/MISO and MIMO arguments is
According to the quadratic approximation, introducing the MIMO scheme should reduce the $x_{\mathrm{3dB}}$ argument by a factor $\sqrt{2}$, which corresponds to $\frac{\alpha_{\mathrm{SIMO/MISO}}}{\alpha_{\mathrm{MIMO}}}$ equal to $\sqrt{2}$. Notice that the parameter $c$, which controls the rate of the decay of the AF mainlobe, does not affect the ratio of the arguments. This confirms the observation from Table I that, although the rate of decay varies across array geometries, the improvement in the argument offered by MIMO is identical. The worst-case relative error introduced by the approximation is below $2.33\%$, matching well with the numerical results. Fig. 1 illustrates the ambiguity function for each considered array geometry and processing. As anticipated, MIMO processing improves the resolution by offering an array factor that decays at a faster rate. Note that the nulls of the ambiguity function remain the same regardless of the processing, as squaring any arbitrary real function preserves the locations of the zeros.
Moreover, the MIMO processing significantly improves the peak-to-sidelobe level. Squaring of the AF corresponds to multiplication of the decay rate by 2, resulting in a twofold improvement of the PSL, as can be seen in Fig. 1. Table II presents the PSL per antenna geometry and processing. Interestingly, the arrays distributed on a plane (URA, UPCA) offer much better PSL than the one-dimensional arrays (ULA, UCA), although the latter offer better resolution.
TABLE II: Peak-to-sidelobe level per array geometry and processing.
Array geometry $\mathrm{PSL_{SIMO/MISO}}$ [dB] $\mathrm{PSL_{MIMO}}$ [dB]
ULA -8.78 -17.57
UCA -7.9 -17.8
URA -17.57 -35.14
UPCA -13.26 -26.52
@@
Figure 1: Ambiguity function for selected antenna arrays and processing. The array aperture is 4m with a center frequency of 6 GHz, the target is located at 10m from the array.

### v2

# Analysis of the Ambiguity Function of Narrowband Near-field MIMO Radar

@@ Abstract
This paper compares the sensing performance of a narrowband near-field system across several practical antenna array geometries and SIMO/MISO and MIMO configurations. For identical transmit and receive apertures, MIMO processing is equivalent to squaring the near-field array factor, resulting in improved beamdepth and sidelobe level. Analytical derivations, supported by simulations, show that the MIMO processing improves the maximum near-field sensing range and resolution by approximately a factor of 1.4 compared to a single-aperture system. Using a quadratic approximation of the mainlobe of the array factor, an analytical improvement factor of $\sqrt{2}$ is derived, validating the numerical results. Finally, MIMO is shown to improve the poor sidelobe performance observed in the near-field by a factor of two, due to squaring of the array factor.
@@ I-A Problem Statement
The radiative near-field (NF) systems are regarded as a key enabler of the next generation of wireless systems [1]. Their ability to create beamspots in the distance domain enables user multiplexing both in range and angle [2, 3]. This unique property has also attracted the attention of the sensing community as beamfocusing in the range domain can be utilized to localize the targets with limited bandwidth [4, 5]. Despite these advantages, near-field sensing performance is subject to significant shortcomings, such as limited resolution, short maximum near-field range and low peak-to-sidelobe level (PSL). The recent research regarding near-field sensing has primarily focused on single-array geometry and single-aperture systems [6, 3, 5, 7, 4]. To improve the near-field sensing, the performance gains from MIMO sensing have to be identified and quantified for different array geometries.
@@ I-B Relevant works
The benefits of MIMO processing for radar and a far-field scenario were investigated in [8]. In [6], the near-field beam focusing was derived for a uniform square array, based on electromagnetic principles. Generalizations of the array factor (AF) in the near-field for arbitrary uniform rectangular array (URA) and uniform planar circular array (UPCA) are provided in [3]. This work studies the beamfocusing property from a communication-centric perspective. The near-field communications performance and AF of a uniform circular array (UCA) were discussed in [7]. In [4], the near-field sensing resolution was demonstrated using a testbed equipped with a uniform linear antenna array (ULA). An approximation of the near-field range resolution for the same array architecture and SIMO/MISO case is provided in [5]. The sensing performance of a near-field system for point and extended targets was investigated in [9]. This work discusses the NF performance of the SIMO/MISO and MIMO systems regarding the Cramér–Rao lower bound, but does not directly discuss the gains from MIMO for point targets.
@@ I-C Contributions
This work evaluates and compares the narrowband NF sensing performance across several array geometries: ULA, UCA, URA and UPCA. The sensing performance is studied for two cases: SIMO/MISO - single aperture and MIMO - two identical collocated apertures. The gains from employing MIMO are assessed in terms of resolution, maximum NF range and PSL of the ambiguity function. Based on a quadratic approximation, the analytical improvement factor is derived, which matches well with the numerical simulations.
@@ II-A Ambiguity function of near-field sensing system
Consider a collocated narrowband antenna array with $M$ transmitters and $N$ receivers. The location of the of the $m$th transmitter is given by $\mathbf{p}_{m}$ and $n$th receiver by $\mathbf{p}_{n}$. The antenna elements are assumed to be isotropic and mutual coupling is considered negligible. The distance between each antenna element and some point in space is $d_{m}(\mathbf{p})=\left|\left|\mathbf{p}-\mathbf{p}_{m}\right|\right|_{2}$ for transmitters and $d_{n}(\mathbf{p})=\left|\left|\mathbf{p}-\mathbf{p}_{n}\right|\right|_{2}$ for receivers. The bistatic distance is denoted as $d_{m,n}(\mathbf{p})=d_{m}(\mathbf{p})+d_{n}(\mathbf{p})$. The region of interest is located within the radiative near-field of the array and the propagation is modelled as line-of-sight. The channel can be written as
@@
$\Delta d_{m}(\mathbf{p}^{\prime},\mathbf{p})$ $=\Delta d_{m=0}(\mathbf{p}^{\prime},\mathbf{p})+\Delta\delta d_{m}(\mathbf{p}^{\prime},\mathbf{p})$ (4)
$\Delta d_{n}(\mathbf{p}^{\prime},\mathbf{p})$ $=\Delta d_{n=0}(\mathbf{p}^{\prime},\mathbf{p})+\Delta\delta d_{n}(\mathbf{p}^{\prime},\mathbf{p}).$ (5)
The matched filter can be expressed in terms of array factors as
$\mathcal{A}(\mathbf{p}^{\prime},\mathbf{p})=e^{-j2\pi\frac{c}{f}\Delta d_{m=0}(\mathbf{p}^{\prime},\mathbf{p})}\underbrace{\frac{1}{\sqrt{M}}\sum_{m\in M}e^{-j2\pi\frac{f}{c}\Delta\delta d_{m}(\mathbf{p}^{\prime},\mathbf{p})}}_{\mathrm{AF_{TX}}(\mathbf{p}^{\prime},\mathbf{p})}$
×$\qquad\times e^{-j2\pi\frac{c}{f}\Delta d_{n=0}(\mathbf{p}^{\prime},\mathbf{p})}\underbrace{\frac{1}{\sqrt{N}}\sum_{n\in N}e^{-j2\pi\frac{f}{c}\Delta\delta d_{n}(\mathbf{p}^{\prime},\mathbf{p})}}_{\mathrm{AF_{RX}}(\mathbf{p}^{\prime},\mathbf{p})}.$ (6)
@@
Similarly, as in the far-field [8], the MIMO configuration improves the ambiguity function by introducing a multiplication by a second array factor compared to the SIMO/MISO case, where only a single array factor defines the ambiguity function. In the near-field region, the AF depends both on angle and distance from the array, unlike in the far-field, where it is only angle dependent. The distance sensitivity results in an array gain that varies with distance and facilitates the beamfocusing. The distance dependence of the AF enables new applications, allowing the discrimination of targets and users located at the same angle but at different distances without bandwidth, which is impossible in the far field.
@@ II-B Near-field 3 dB beamdepth
Consider a target and points of interest located along the array broadside, which allows for reducing the vectors $\mathbf{p}$ and $\mathbf{p}^{\prime}$ to scalars $d$ and $d^{\prime}$, which represent the distance from the geometric center of the array along the normal. The AF in the near-field can be calculated analytically for different array architectures. [6, 3, 7]. Due to the Fresnel approximation of distance in the NF of the array, the argument of the AF shares the same formula for different array architectures
$x(d,d^{\prime})$ $=ad_{\mathrm{FA}}d_{\mathrm{ver}}$ (8)
where $a$ is an argument scaling coefficient dependent on the array geometry, $d_{\mathrm{FA}}=2D^{2}/\lambda$ is the Fraunhofer distance of the array, $D$ is the array aperture, $\lambda$ is the wavelength and $d_{\mathrm{ver}}=\left|\frac{1}{d}-\frac{1}{d^{\prime}}\right|=\frac{|d^{\prime}-d|}{d^{\prime}d}$ is the absolute value of vergence difference [10]. For brevity in the following, the shorthand $x$ is used to represent $x(d,d^{\prime})$. Consider some near-field ambiguity function given by $|\mathcal{A}(x)|^{2}$. For a SISO/SIMO system, the ambiguity function will be defined by a single array factor of the transmit or receive aperture $|\mathcal{A}(x)|_{\mathrm{SIMO/MISO}}^{2}=|\mathrm{AF}(x)|^{2}$. For the MIMO scheme, the ambiguity function will be a product of two array factors. For simplicity, the transmit and receive arrays are considered to be identical and exactly collocated, resulting in a monostatic system $|\mathcal{A}(x)|_{\mathrm{MIMO}}^{2}=|\mathrm{AF}(x)|^{4}$.
The range resolution is evaluated in terms of the radial half-power beamwidth of the ambiguity function, also referred to as the beamdepth (BD). Note that in the considered narrowband system, the primary source of resolution is the distance-dependent beampattern. The resolution due to bandwidth is considered to be negligibly small compared to the resolution due to beamfocusing. The 3 dB beamdepth is calculated by computing the argument for which the ambiguity function reaches half of the maximum value. Considering normalized ambiguity function it can be written as $|\mathcal{A}(x_{\mathrm{3dB}})|^{2}=0.5$. The argument value is then used to calculate the distance $d$ at which the $-3$ dB value is reached by solving (8) for $d$ resulting in
@@
Formula (9) gives two distances for which the $-3$ dB value is achieved for the target at $d^{\prime}$. Where $\alpha=x_{\mathrm{3dB}}/a$ represents the value of the argument $d_{\mathrm{FA}}d_{\mathrm{ver}}$ for which the normalized AF reaches half of the maximum value. The parameter $a$ depends only on the array geometry, while $x_{\mathrm{3dB}}$ depends on both the array architecture and configuration, SIMO/MISO or MIMO.
As $d_{\mathrm{3dB}}$ should always be positive, the following constraint arises $d^{\prime}<\frac{d_{\mathrm{FA}}}{\alpha}$. Given the $-3$ dB crossing points, 3 dB beamdepth (BD) is
$\mathrm{BD}$ =$=\begin{cases}\frac{2\alpha d_{\mathrm{FA}}d^{\prime 2}}{d_{\mathrm{FA}}^{2}-\alpha^{2}d^{\prime 2}},&d^{\prime}<\frac{d_{\mathrm{FA}}}{\alpha},\\ \infty,&d^{\prime}\geq\frac{d_{\mathrm{FA}}}{\alpha}.\end{cases}$ (10)
For $d$ greater or equal to $\frac{d_{\mathrm{FA}}}{\alpha}$ the beamdepth extends from $d_{\mathrm{FA}}/\alpha$ to infinity. In the context of beamfocusing, the maximum near-field range is defined by a maximum distance for which the beamdepth remains finite, which is $d_{\mathrm{FA}}/\alpha$. Note that parameter $\alpha$ determines both the beamdepth and the maximum beamfocusing range, with smaller alpha corresponding to better performance. In the following sections, parameter $\alpha$ is computed for different array architectures to assess the resolution and maximum near-field sensing range. The target is assumed to be located on the array broadside, which corresponds to the best-case scenario offering the highest resolution due to the largest effective aperture. For other angles, the effective aperture size is decreased and can be approximated as a projection of the array on a plane that is perpendicular to the axis of the target [3].
@@ III-D2 MIMO
ULA 6.952 4.969 1.399
UCA 5.737 4.148 1.383
URA 9.937 7.068 1.406
UPCA 7.087 5.103 1.389
@@
Figure 1: Ambiguity function for selected antenna arrays and processing. The array aperture is 4m with a center frequency of 6 GHz and the target is 10m from the array.
Table I presents the summarized values of the parameter $\alpha$ for each considered array geometry and processing. By comparing the $\alpha$ values across antenna geometries and processing a few observations can be made. For the same array aperture, UCA offers the best resolution. The MIMO processing improves the parameter $\alpha$ approximately by the same factor of $1.4$ regardless of the array geometry. This corresponds to $1.4$-fold improvement in the NF sensing resolution and maximum NF range.
The consistent ratio $\frac{\alpha_{\mathrm{SIMO/MISO}}}{\alpha_{\mathrm{MIMO}}}$ observed across various array geometries can be attributed to quadratic decay of AF within the mainlobe. Based on this, the mainlobe of the AF can be accurately approximated as a quadratic function [11]. For small arguments, the NF AF mainlobe is an even function, which for SIMO/MISO can be approximated as
$|\mathrm{AF}(x)|^{2}$ $\approx 1-cx^{2},$ (23)
where $c$ is some scalar that controls the decay rate and depends on the array geometry. Similarly, as in the previous subsections, the $-3$ dB value is obtained for the argument $x_{\mathrm{3dB}}^{\mathrm{SIMO/MISO}}=\frac{\sqrt{2}}{2\sqrt{c}}$. In the MIMO case, the approximation from (23) is squared. To keep the approximation quadratic for both scenarios, higher order terms are neglected, resulting in
$|\mathrm{AF}(x)|^{4}$ $\approx 1-2cx^{2}.$ (24)
The $-3$ dB value is obtained for argument $x_{\mathrm{3dB}}^{\mathrm{MIMO}}=\frac{1}{2\sqrt{c}}$. Consequently, the ratio between the SIMO/MISO and MIMO arguments is
According to the quadratic approximation, introducing the MIMO scheme should reduce the $x_{\mathrm{3dB}}$ argument by a factor $\sqrt{2}$, which corresponds to $\frac{\alpha_{\mathrm{SIMO/MISO}}}{\alpha_{\mathrm{MIMO}}}$ equal to $\sqrt{2}$. Notice that the parameter $c$, which controls the rate of the decay of the AF mainlobe, does not affect the ratio of the arguments. This confirms the observation from Table I that, although the rate of decay varies across array geometries, the improvement in the argument offered by MIMO is identical. The worst-case relative error introduced by the approximation is below $2.27\%$, matching well with the numerical results. Fig. 1 illustrates the ambiguity function for each considered array geometry and processing. As anticipated, MIMO processing improves the resolution by offering an array factor that decays at a faster rate. Note that the nulls of the ambiguity function remain the same regardless of the processing, as squaring any arbitrary real function preserves the locations of the zeros.
Moreover, the MIMO processing significantly improves the peak-to-sidelobe level. Squaring of the AF corresponds to multiplication of the decay rate by 2, resulting in a twofold improvement of the PSL, as can be seen in Fig. 1. Table II presents the PSL per antenna geometry and processing. Interestingly, the arrays distributed on a plane (URA, UPCA) offer much better PSL than the one-dimensional arrays (ULA, UCA), although the latter provide better resolution.
TABLE II: Peak-to-sidelobe level per array geometry and processing.
Array geometry $\mathrm{PSL_{SIMO/MISO}}$ [dB] $\mathrm{PSL_{MIMO}}$ [dB]
ULA -8.78 -17.57
UCA -7.90 -15.80
URA -17.57 -35.13
UPCA -13.26 -26.52